# Stochastic Monkeys at Play: Random Augmentations Cheaply Break LLM Safety Alignment

### v1

@@ Appendix C Expanded Experimental Results
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.191 +0.018 +0.018 +0.004 +0.013 +0.087 +0.084 +0.131 +0.101 +0.057
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.218 +0.009 +0.011 -0.007 +0.004 +0.082 +0.073 +0.104 +0.087 +0.046
@@
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.271 +0.029 +0.009 +0.044 +0.027 +0.116 +0.107 +0.200 +0.141 +0.084
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.289 +0.044 +0.009 +0.044 +0.033 +0.149 +0.122 +0.200 +0.157 +0.095
@@
0.7 0 0.271 +0.029 +0.009 +0.136 +0.058 +0.282 +0.182 +0.258 +0.241 +0.149
1.0 0 0.289 +0.044 +0.009 +0.144 +0.066 +0.298 +0.189 +0.293 +0.260 +0.163
Llama 3.1 8B Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.269 +0.067 +0.009 +0.047 +0.041 +0.136 +0.129 +0.218 +0.161 +0.101
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.311 +0.038 -0.020 +0.016 +0.011 +0.104 +0.084 +0.184 +0.124 +0.068
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.309 +0.033 -0.020 -0.004 +0.003 +0.082 +0.071 +0.173 +0.109 +0.056
0.0 0 0.269 +0.067 +0.009 +0.138 +0.071 +0.278 +0.184 +0.309 +0.257 +0.164
0.7 0 0.311 +0.038 -0.020 +0.098 +0.039 +0.249 +0.138 +0.273 +0.220 +0.129
1.0 0 0.309 +0.033 -0.020 +0.098 +0.037 +0.229 +0.138 +0.260 +0.209 +0.123
@@
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.727 +0.156 +0.107 +0.107 +0.123 +0.144 +0.156 +0.160 +0.153 +0.138
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.740 +0.158 +0.104 +0.111 +0.124 +0.169 +0.164 +0.164 +0.166 +0.145
@@
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.202 +0.053 +0.060 +0.073 +0.062 +0.187 +0.173 +0.207 +0.189 +0.126
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.229 +0.071 +0.096 +0.069 +0.079 +0.202 +0.187 +0.238 +0.209 +0.144
@@
0.7 0 0.202 +0.053 +0.060 +0.158 +0.090 +0.376 +0.271 +0.318 +0.321 +0.206
1.0 0 0.229 +0.071 +0.096 +0.200 +0.122 +0.422 +0.324 +0.367 +0.371 +0.247
Qwen 2 7B Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.356 +0.109 +0.116 +0.098 +0.107 +0.233 +0.213 +0.262 +0.236 +0.172
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.396 +0.124 +0.142 +0.124 +0.130 +0.222 +0.193 +0.240 +0.219 +0.174
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.438 +0.147 +0.162 +0.116 +0.141 +0.220 +0.198 +0.253 +0.224 +0.183
0.0 0 0.356 +0.109 +0.116 +0.207 +0.144 +0.360 +0.269 +0.336 +0.321 +0.233
0.7 0 0.396 +0.124 +0.142 +0.227 +0.164 +0.360 +0.256 +0.333 +0.316 +0.240
1.0 0 0.438 +0.147 +0.162 +0.218 +0.176 +0.367 +0.278 +0.329 +0.324 +0.250
Vicuna 7B v1.5 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.369 +0.096 +0.096 +0.113 +0.101 +0.198 +0.184 +0.196 +0.193 +0.147
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.416 +0.080 +0.084 +0.084 +0.083 +0.131 +0.142 +0.173 +0.149 +0.116
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.502 +0.073 +0.076 +0.071 +0.073 +0.104 +0.140 +0.142 +0.129 +0.101
0.0 0 0.369 +0.096 +0.096 +0.209 +0.133 +0.304 +0.251 +0.264 +0.273 +0.203
0.7 0 0.416 +0.080 +0.084 +0.180 +0.115 +0.280 +0.224 +0.253 +0.253 +0.184
1.0 0 0.502 +0.073 +0.076 +0.191 +0.113 +0.269 +0.216 +0.236 +0.240 +0.177
@@
0.7 0 0.884 +0.064 +0.073 +0.089 +0.076 +0.089 +0.100 +0.087 +0.092 +0.084
1.0 0 0.882 +0.064 +0.076 +0.087 +0.076 +0.098 +0.100 +0.096 +0.098 +0.087
Table 4: $(25,\gamma)$-success rate gains of models $f^{\prime}$ over the smallest model in their model family $f$, using greedy decoding for $g$. Adjacent rows are grouped by model family. An asterisk (*) next to a model name indicates the model $f$ is the smallest in its respective model family, and the values for that row report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$.
@@
Qwen 2 0.5B Instruct* $\gamma^{*}_{\mathcal{A}}$ 0.431 0.578 0.633 0.598 0.603 0.731 0.767 0.756 0.751 0.677
Qwen 2 1.5B Instruct $\gamma^{*}_{\mathcal{A}}$0.111 -0.196 -0.249 -0.204 -0.216 -0.213 -0.262 -0.218 -0.231 -0.224
Qwen 2 7B Instruct $\gamma^{*}_{\mathcal{A}}$0.076 -0.113 -0.162 -0.144 -0.140 -0.142 -0.198 -0.138 -0.159 -0.150
Qwen 2 0.5B Instruct* 0 0.431 0.578 0.633 0.713 0.641 0.867 0.867 0.844 0.859 0.750
Qwen 2 1.5B Instruct 0 -0.111 -0.196 -0.249 -0.216 -0.220 -0.202 -0.278 -0.224 -0.235 -0.227
Qwen 2 7B Instruct 0 -0.076 -0.113 -0.162 -0.151 -0.142 -0.151 -0.242 -0.153 -0.182 -0.162
Vicuna 7B v1.5* $\gamma^{*}_{\mathcal{A}}$ 0.369 0.464 0.464 0.482 0.470 0.567 0.553 0.564 0.561 0.516
Vicuna 13B v.15 $\gamma^{*}_{\mathcal{A}}$0.082 -0.118 -0.102 -0.127 -0.116 -0.122 -0.129 -0.107 -0.119 -0.117
Vicuna 7B v1.5* 0 0.369 0.464 0.464 0.578 0.502 0.673 0.620 0.633 0.642 0.572
Vicuna 13B v.15 0 -0.082 -0.118 -0.102 -0.124 -0.115 -0.111 -0.118 -0.093 -0.107 -0.111
Table 5: $(25,\gamma)$-success rate gains of quantized models $f^{\prime}$ over their respective original model $f$, using greedy decoding for $g$. Base model rows are indicated with “None” and report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$.
@@
Llama 3.1 8B Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.269 0.336 0.278 0.316 0.310 0.404 0.398 0.487 0.430 0.370
W8A8 $\gamma^{*}_{\mathcal{A}}$0.027 -0.024 -0.009 -0.020 -0.018 -0.036 -0.031 -0.024 -0.030 -0.024
W4A16 $\gamma^{*}_{\mathcal{A}}$0.042 +$0$0.022 -0.024 -0.016 -0.031 -0.047 -0.029 -0.036 -0.026
None 0 0.269 0.336 0.278 0.407 0.340 0.547 0.453 0.578 0.526 0.433
W8A8 0 -0.027 -0.024 -0.009 -0.018 -0.017 -0.011 -0.022 -0.007 -0.013 -0.015
W4A16 0 -0.042 +$0$0.022 -0.009 -0.010 -0.013 -0.033 -0.029 -0.025 -0.018
@@
Qwen 2 7B Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.356 0.464 0.471 0.453 0.463 0.589 0.569 0.618 0.592 0.527
W8A8 $\gamma^{*}_{\mathcal{A}}$ +0.020 +0.036 +0.044 +0.042 +0.041 +0.029 +0.016 +0.011 +0.019 +0.030
W4A16 $\gamma^{*}_{\mathcal{A}}$0.178 -0.076 -0.091 -0.144 -0.104 -0.231 -0.164 -0.178 -0.191 -0.147
None 0 0.356 0.464 0.471 0.562 0.499 0.716 0.624 0.691 0.677 0.588
W8A8 0 +0.020 +0.036 +0.044 +0.038 +0.039 +0.027 +0.016 +0.013 +0.019 +0.029
W4A16 0 -0.178 -0.076 -0.091 -0.102 -0.090 -0.180 -0.133 -0.140 -0.151 -0.120
Vicuna 7B v1.5 None $\gamma^{*}_{\mathcal{A}}$ 0.369 0.464 0.464 0.482 0.470 0.567 0.553 0.564 0.561 0.516
W8A8 $\gamma^{*}_{\mathcal{A}}$0.020 -0.007 +0.004 -0.031 -0.011 -0.027 -0.016 -0.002 -0.015 -0.013
W4A16 $\gamma^{*}_{\mathcal{A}}$ +0.020 +0.058 +0.053 +0.049 +0.053 +0.060 +0.069 +0.080 +0.070 +0.061
None 0 0.369 0.464 0.464 0.578 0.502 0.673 0.620 0.633 0.642 0.572
W8A8 0 -0.020 -0.007 +0.004 -0.013 -0.005 -0.002 -0.004 -0.002 -0.003 -0.004
W4A16 0 +0.020 +0.058 +0.053 +0.071 +0.061 +0.078 +0.073 +0.084 +0.079 +0.070
@@
Table 6: $(25,\gamma)$-success rate gains of models $f^{\prime}$ with fine-tuning-based defenses over their respective base models $f$, using greedy decoding for $g$. Adjacent rows are grouped into pairs of the base model and its fine-tuned version. Base model rows report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$.
@@
Appendix D Example Model Responses
Figures 7-13 show example model responses to harmful prompts under random augmentations.
@@ Appendix D Example Model Responses
Figure 12: String insertion (suffix) on a SORRY-Bench prompt for W8A8 quantized Qwen 2 7B Instruct. Both the original model (“FP16”) and the quantized model refuse to comply with the original prompt. Applying the same set of random augmentations to each model reveals an interesting discrepancy: most responses for both models are refusals, but for the augmentation shown a the original model refuses whereas the W8A8 model complies.

### v2

@@ D.4.2 Results
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.236 -0.027 -0.027 -0.040 -0.031 +0.042 +0.040 +0.087 +0.056 +0.013
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.260 -0.033 -0.031 -0.049 -0.038 +0.040 +0.031 +0.062 +0.044 +0.003
@@
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.387 -0.087 -0.107 -0.071 -0.088 +$0$0.009 +0.084 +0.025 -0.031
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.449 -0.116 -0.151 -0.116 -0.127 -0.011 -0.038 +0.040 -0.003 -0.065
@@
0.7 0 0.387 -0.087 -0.107 +0.020 -0.058 +0.167 +0.067 +0.142 +0.125 +0.034
1.0 0 0.449 -0.116 -0.151 -0.016 -0.094 +0.138 +0.029 +0.133 +0.100 +0.003
Llama 3.1 8B Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.140 +0.029 +0.060 +0.024 +0.038 +0.053 +0.042 +0.116 +0.070 +0.054
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.236 -0.067 -0.027 -0.071 -0.055 -0.056 -0.051 +0.007 -0.033 -0.044
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.340 -0.171 -0.136 -0.180 -0.162 -0.160 -0.156 -0.087 -0.134 -0.148
0.0 0 0.140 +0.029 +0.060 +0.073 +0.054 +0.142 +0.084 +0.191 +0.139 +0.097
0.7 0 0.236 -0.067 -0.027 -0.016 -0.036 +0.058 -0.013 +0.089 +0.044 +0.004
1.0 0 0.340 -0.171 -0.136 -0.107 -0.138 -0.069 -0.109 -0.011 -0.063 -0.100
@@
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.893 -0.011 -0.060 -0.060 -0.044 -0.022 -0.011 -0.007 -0.013 -0.029
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.916 -0.018 -0.071 -0.064 -0.051 -0.007 -0.011 -0.011 -0.010 -0.030
@@
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.333 -0.078 -0.071 -0.058 -0.069 +0.056 +0.042 +0.076 +0.058 -0.006
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.400 -0.100 -0.076 -0.102 -0.093 +0.031 +0.016 +0.067 +0.038 -0.027
@@
0.7 0 0.333 -0.078 -0.071 +0.027 -0.041 +0.244 +0.140 +0.187 +0.190 +0.075
1.0 0 0.400 -0.100 -0.076 +0.029 -0.049 +0.251 +0.153 +0.196 +0.200 +0.076
Qwen 2 7B Instruct 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.467 +0.100 +0.091 +0.102 +0.098 +0.216 +0.160 +0.198 +0.191 +0.144
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.716 -0.107 -0.107 -0.104 -0.106 -0.022 -0.060 -0.020 -0.034 -0.070
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.773 -0.113 -0.120 -0.131 -0.121 -0.031 -0.047 -0.027 -0.035 -0.078
0.0 0 0.467 +0.100 +0.091 +0.198 +0.130 +0.329 +0.236 +0.293 +0.286 +0.208
0.7 0 0.716 -0.107 -0.107 +$0$0.071 +0.116 +0.016 +0.062 +0.064 -0.003
1.0 0 0.773 -0.113 -0.120 -0.018 -0.084 +0.071 +0.018 +0.038 +0.042 -0.021
Vicuna 7B v1.5 0.0 $\gamma^{*}_{\mathcal{A}}$ 0.413 +0.100 +0.098 +0.102 +0.100 +0.182 +0.176 +0.200 +0.186 +0.143
0.7 $\gamma^{*}_{\mathcal{A}}$ 0.767 -0.211 -0.224 -0.218 -0.218 -0.147 -0.156 -0.118 -0.140 -0.179
1.0 $\gamma^{*}_{\mathcal{A}}$ 0.873 -0.242 -0.240 -0.260 -0.247 -0.200 -0.191 -0.173 -0.188 -0.218
0.0 0 0.413 +0.100 +0.098 +0.191 +0.130 +0.311 +0.244 +0.267 +0.274 +0.202
0.7 0 0.767 -0.211 -0.224 -0.120 -0.185 -0.013 -0.080 -0.040 -0.044 -0.115
1.0 0 0.873 -0.242 -0.240 -0.133 -0.205 -0.047 -0.118 -0.102 -0.089 -0.147
@@
0.7 0 0.971 -0.022 -0.013 +0.002 -0.011 +0.002 +0.013 +0.000 +0.005 -0.003
1.0 0 0.980 -0.033 -0.022 -0.011 -0.022 +0.000 +0.002 -0.002 +$0$0.011
Figure 12: Average $\left(25,\gamma^{*}_{\mathcal{A}}\right)$-success rate gains of different kinds of augmentations over using no augmentations, using temperature sampling $\tau\in\{0.7,1.0\}$ for generation.
Table 8: $(25,\gamma)$-success rate gains of models $f^{\prime}$ over the smallest model in their model family $f$, using greedy decoding for $g$. Adjacent rows are grouped by model family. An asterisk (*) next to a model name indicates the model $f$ is the smallest in its respective model family, and the values for that row report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The largest absolute value among string insertion augmentations and among character-level augmentations is bolded. Additionally, the average for both kinds of augmentations is reported, and the larger absolute value is bolded. The rightmost column reports the overall average over both kinds of augmentations.
@@
Qwen 2 0.5B Instruct* $\gamma^{*}_{\mathcal{A}}$ 0.480 0.649 0.676 0.627 0.650 0.760 0.800 0.771 0.777 0.714
Qwen 2 1.5B Instruct $\gamma^{*}_{\mathcal{A}}$0.138 -0.244 -0.249 -0.211 -0.235 -0.224 -0.267 -0.196 -0.229 -0.232
Qwen 2 7B Instruct $\gamma^{*}_{\mathcal{A}}$0.013 -0.082 -0.118 -0.058 -0.086 -0.078 -0.173 -0.107 -0.119 -0.103
Qwen 2 0.5B Instruct* 0 0.480 0.649 0.676 0.747 0.690 0.878 0.851 0.836 0.855 0.773
Qwen 2 1.5B Instruct 0 -0.138 -0.244 -0.249 -0.213 -0.236 -0.180 -0.238 -0.176 -0.198 -0.217
Qwen 2 7B Instruct 0 -0.013 -0.082 -0.118 -0.082 -0.094 -0.082 -0.149 -0.076 -0.102 -0.098
Vicuna 7B v1.5* $\gamma^{*}_{\mathcal{A}}$ 0.413 0.513 0.511 0.516 0.513 0.596 0.589 0.613 0.599 0.556
Vicuna 13B v.15 $\gamma^{*}_{\mathcal{A}}$0.093 -0.109 -0.104 -0.120 -0.111 -0.062 -0.089 -0.080 -0.077 -0.094
Vicuna 7B v1.5* 0 0.413 0.513 0.511 0.604 0.543 0.724 0.658 0.680 0.687 0.615
Vicuna 13B v.15 0 -0.093 -0.109 -0.104 -0.091 -0.101 -0.080 -0.067 -0.078 -0.075 -0.088
Table 9: $(25,\gamma)$-success rate gains of quantized models $f^{\prime}$ over their respective original model $f$, using greedy decoding for $g$. Base model rows are indicated with “None” and report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The largest absolute value among string insertion augmentations and among character-level augmentations is bolded. Additionally, the average for both kinds of augmentations is reported, and the larger absolute value is bolded. The rightmost column reports the overall average over both kinds of augmentations.
@@
Llama 3.1 8B Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.140 0.169 0.200 0.164 0.178 0.193 0.182 0.256 0.210 0.194
W8A8 $\gamma^{*}_{\mathcal{A}}$0.004 +0.002 +0.004 -0.004 +0.001 -0.007 -0.011 +0.007 -0.004 -0.001
W4A16 $\gamma^{*}_{\mathcal{A}}$ +0.040 +0.073 +0.031 +0.047 +0.050 +0.042 +0.069 +0.067 +0.059 +0.055
None 0 0.140 0.169 0.200 0.213 0.194 0.282 0.224 0.331 0.279 0.237
W8A8 0 -0.004 +0.002 +0.004 +0.004 +0.004 +0.018 +0.009 +0.002 +0.010 +0.007
W4A16 0 +0.040 +0.073 +0.031 +0.113 +0.073 +0.098 +0.096 +0.084 +0.093 +0.083
@@
Qwen 2 7B Instruct None $\gamma^{*}_{\mathcal{A}}$ 0.467 0.567 0.558 0.569 0.564 0.682 0.627 0.664 0.658 0.611
W8A8 $\gamma^{*}_{\mathcal{A}}$ +0.007 +0.033 +0.038 +0.011 +0.027 +0.000 +0.016 +0.027 +0.014 +0.021
W4A16 $\gamma^{*}_{\mathcal{A}}$0.251 -0.087 -0.060 -0.147 -0.098 -0.216 -0.122 -0.151 -0.163 -0.130
None 0 0.467 0.567 0.558 0.664 0.596 0.796 0.702 0.760 0.753 0.674
W8A8 0 +0.007 +0.033 +0.038 +0.013 +0.028 +0.013 +0.004 +0.000 +0.006 +0.017
W4A16 0 -0.251 -0.087 -0.060 -0.096 -0.081 -0.142 -0.098 -0.160 -0.133 -0.107
Vicuna 7B v1.5 None $\gamma^{*}_{\mathcal{A}}$ 0.413 0.513 0.511 0.516 0.513 0.596 0.589 0.613 0.599 0.556
W8A8 $\gamma^{*}_{\mathcal{A}}$0.040 -0.018 -0.024 -0.011 -0.018 -0.027 -0.011 -0.029 -0.022 -0.020
W4A16 $\gamma^{*}_{\mathcal{A}}$0.002 +0.082 +0.051 +0.060 +0.064 +0.073 +0.093 +0.087 +0.084 +0.074
None 0 0.413 0.513 0.511 0.604 0.543 0.724 0.658 0.680 0.687 0.615
W8A8 0 -0.040 -0.018 -0.024 -0.002 -0.015 +0.004 -0.009 -0.020 -0.008 -0.011
W4A16 0 -0.002 +0.082 +0.051 +0.082 +0.072 +0.080 +0.084 +0.080 +0.081 +0.077
@@
Table 10: $(25,\gamma)$-success rate gains of models $f^{\prime}$ with fine-tuning-based defenses over their respective base models $f$, using greedy decoding for $g$. Adjacent rows are grouped into pairs of the base model and its fine-tuned version. Base model rows report the empirical $(25,\gamma)$-success rate $\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The other rows report the empirical $(25,\gamma)$-success rate gain $\hat{r}_{25,\gamma}(\mathcal{A},f^{\prime},g)-\hat{r}_{25,\gamma}(\mathcal{A},f,g)$. The largest absolute value among string insertion augmentations and among character-level augmentations is bolded. Additionally, the average for both kinds of augmentations is reported, and the larger absolute value is bolded. The rightmost column reports the overall average over both kinds of augmentations.
@@
D.5 Example Model Responses
Figures 13-19 show example model responses to harmful prompts under random augmentations.
@@ D.5 Example Model Responses
Figure 18: String insertion (suffix) on a SORRY-Bench prompt for W8A8 quantized Qwen 2 7B Instruct. Both the original model (“FP16”) and the quantized model refuse to comply with the original prompt. Applying the same set of random augmentations to each model reveals an interesting discrepancy: most responses for both models are refusals, but for the final augmentation shown the original model refuses whereas the W8A8 model complies.